OpenStudy (kamille):

Hello, I need to solve this: \[5^{\log _{0,7}(\log _{2}(\frac{ 3+5x }{ x ^{2} +3})}<1\]

OpenStudy (kamille):

@kausarsalley that's is the problem

OpenStudy (anonymous):

okay...i see it....i am on it

OpenStudy (kamille):

what have I done: \[5^{\log _{0,7}(\log _{2}(\frac{ 3+5x }{ x ^{2} +3})}<5^{0}\]

OpenStudy (anonymous):

wait, is that the question because i cannot see it clearly and I want to be sure